A tank contains water on top of mercury.A cube of iron,60mm along each edge,is sitting upright in equilibrium in the liquids.Find how much of it is in each liquid.The densities of iron and mercury are 7.7g/cm3 and 13.6g/cm3 respectively

Grade:12

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Sohan

Suppose x1 length is in water and x2 in mercury

and x1+x2=h=60mm...............(1)

here equilibrium equation will be

Weight = total buoyant force for each medium,  W=B=Bw + Bhg...........(2)

weight=W= density of cubex volumexg= dcx(Axh)xg,   where A is h2

Bw=dw*(A*x1)*g

Bhg=dhg*(A*x2)*g

use equation 2

after cancelling terms on rhs and lhs you will get

dc*h=dw*x1+dhg*x2

put values of density

7.7h=x1+13.6x2..............(3)

solve for x1 and x2 using 1 and 3

x2=6.7h/12.6= 6.7*60/12.6 = 31.9 mm in mercury

in water=60-31.9=28.1 mm

 All the best.                                                          

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..